# A Survey on All-in-One Image Restoration: Taxonomy, Evaluation and Future Trends

### v2

@@ III-B5 Machine Unlearning
Privacy protection is an increasingly critical issue in the field of artificial intelligence, particularly as models become more integrated into sensitive applications. One promising solution to address this challenge is machine unlearning [132]. This innovative approach aims to effectively erase the influence of private data from trained models, allowing them to function as if that sensitive information was never utilized during the training process. Existing methodologies for unlearning can be broadly categorized into two types: exact unlearning [133], which seeks to completely eliminate the impact of specific data points, and approximate unlearning [134], which aims to diminish the influence of such data to a certain degree. While the concept of unlearning has been extensively explored in various contexts, including classification tasks and federated learning scenarios, its application in the realm of end-to-end IR remains largely uncharted.
To fill this gap, Su et al. [135] proposed an innovative framework that applied machine unlearning techniques to AiOIR models. They defined a scenario in which certain types of degradation are treated as private information that must be effectively forgotten by the model. Importantly, this process is designed to preserve the model’s performance on other types of degradations, ensuring that the overall functionality remains intact. To achieve this goal, the authors introduced a technique called instance-wise unlearning, which leveraged adversarial examples in conjunction with gradient ascent methods. This approach not only enhances the model’s ability to forget specific data but also maintains its robustness across various image restoration tasks.
@@ III-C Key Improvements of AiOIR
In addition to the network design in Sec. III-A and learning strategies in Sec. III-B, there are other key techniques to improve AiOIR models. In this section, we preliminarily categorize the key improvements of AiOIR models into the following four areas, i.e., Prompt Learning, Mixture-of-Experts (MoE), Multimodal Model and Agentic IR System, which correspond to Sec. III-C1, Sec. III-C2, Sec. III-C3, Sec. III-C4, respectively. We also illustrate some other improvements, including deep unfolding methods, mask image modeling, hypernetworks, etc. For the sake of clarity, we list the representative works on all-in-one models based on the four types in Fig. 4.
@@ III-C2 Mixture-of-Experts
Yang et al. [99] proposed language-driven all-in-one adverse weather removal. This method generates degradation priors based on textual descriptions of weather conditions, which are then used to guide the adaptive selection of restoration experts. Similarly, WM-MoE [151] introduces a weather-aware routing mechanism to direct image tokens to specialized experts and employs multi-scale experts to handle diverse weather conditions. Both WM-MoE and MoFME [152] enhance IR tasks and downstream tasks like segmentation, while significantly reducing model parameters and inference time. In contrast, MEASNet [82] presents a novel multi-expert adaptive selection mechanism. This mechanism leverages both local and global image features to select the most appropriate expert models, effectively balancing task-specific demands and promoting resource sharing. More recently, MoCE-IR [153] introduces complexity experts with adaptive variable-size computational units that use a spring-inspired force mechanism to dynamically allocate experts. Building on similar principles, UniRestorer [154] construct a multi-granularity degradation set and a multi-granularity MoE restoration model. LoRA-IR [155] leverages a novel mixture of low-rank experts structure, combined with a CLIP-based degradation-guided router.
@@ III-C5 Other Methods
In addition to highlighting four key improvements, we also review other methods for AiOIR. Some models benefit from iterative algorithms in the deep unfolding networks (DUNs), and integrating the network design with pretrained mask image modeling (MIM) also holds significant potential. Meanwhile, hypernetworks with dynamically generated parameters are also highly suitable for AiOIR. The exploration of visual autoregressive modeling (VAR) as a novel paradigm for low-level vision tasks has recently extended to the field of AiOIR. The details are as follows.
Deep Unfolding Methods. Under the framework of optimization theory and deep learning, DUNs enhance IR by unfolding iterative functions into end-to-end trainable architectures. Zhang et al. [164] proposed a DUN framework that integrates CNNs with model-based priors for task-specific IR. Given a degradation model, the clean image is estimated by minimizing an energy function. The half-quadratic splitting (HQS) algorithm [165] decomposes the optimization into two subproblems: one for the data fidelity term and the other for the prior. These subproblems are alternately solved, with the data term handled via least squares and the prior term approximated using CNNs. Building on this principle, DRM-IR [60] enhances the flexibility of all-in-one scenarios by introducing a reference-based and task-adaptive modeling paradigm. It establishes an efficient AiOIR framework that jointly models task-adaptive degradation and model-based restoration. Recently, Up-restore [166] employs a prompt-driven unrolling scheme inspired by ADMM and grounded in MAP estimation, effectively combining data-driven learning with physical priors. VLU-Net [149] leverages an automatic gradient estimation strategy based on VLMs, integrating multi-level information to improve feature preservation.
@@
  Image Deblurring GoPro [167] 2017 2103 $1$ Blurred images at 1280x720 real
@@
KITTI-snow [184] 2023 50 - Two intensity levels: severe and extremely severe synthetic
@@
Mask Image Modeling. Mask image modeling (MIM) (e.g.,  [197, 144, 198]) is a technique in CV that involves training models to predict masked portions of images based on their surrounding context inspired by mask language modeling [199, 200]. This self-supervised approach reconstructs missing image regions, enhancing its visual representation learning. MIM has demonstrated effectiveness in visual tasks such as classification, detection, and segmentation. The MAE [144] framework effectively employs MIM to predict hidden tokens, showcasing impressive performance and generalization capabilities across a range of downstream tasks. Qin et al. [201] are the pioneers in introducing MIM into AiOIR, shifting the focus from degradation‐type discrimination to intrinsic image information recovery. They proposed mask attribute conductance (MAC) to analyze each layer’s contribution to resolving the input integrity gap and rank layers in descending order. DyNet  [202] trains two variants (small and large) concurrently to reconstruct the clean image from masked degraded inputs. Notably, both variants share weights via an intra-network block-repetition scheme, reducing GPU hours by 50%. Most recently, Cat-AIR  [203] introduces a content-aware and task-aware router that dynamically applies self-attention to complex regions and convolution to simpler ones, using learned masks to distinguish hard from easy patches within the feature map.
@@ IV Experiments
To facilitate a comprehensive and effective comparison of various AiOIR methods, we begin by summarizing the key datasets, experimental settings, and evaluation metrics commonly used across different tasks. Following this, we conduct a detailed comparison of existing benchmarks across several representative image restoration tasks, such as low-light enhancement, dehazing, deblurring, denoising, deraining, and desnowing. This structured schema ensures a thorough evaluation of the performance and capabilities of different all-in-one image restoration methods.
@@ IV-A Datasets and Evaluation Metrics
  Single LPN [222] CVPR’19 3M 20.84/0.828 24.88/0.784 26.47/0.778 24.77/0.748 21.26/0.552 23.64/0.738 Specific
@@
Multiple MPRNet [225] CVPR’21 16M 28.00/0.958 33.86/0.958 33.27/0.920 30.76/0.871 27.29/0.761 30.63/0.894 General
@@
All-in-One DL[228] TPAMI’19 2M 26.92/0.931 32.62/0.931 33.05/0.914 30.41/0.861 26.90/0.740 29.98/0.875 parameterized image operator
@@
[…]
@@
VLU-Net [149] $\bigstar$ CVPR’25 35M+88M 30.71/0.980 38.93/0.984 34.13/0.935 31.48/0.892 28.23/0.804 32.70/0.919 vision-language gradient descent-driven
@@ IV-B Experiments and Analysis
  Single ADFNet [223] AAAI’23 8M 24.18/0.928 32.97/0.943 31.15/0.882 25.79/0.781 21.15/0.823 27.05/0.871 Specific
@@
Multiple SwinIR [248] ICCVW’21 1M 21.50/0.891 30.78/0.923 30.59/0.868 24.52/0.773 17.81/0.723 25.04/0.835 General
@@
All-in-One DL [228] TPAMI’19 2M 20.54/0.826 21.96/0.762 23.09/0.745 19.86/0.672 19.83/0.712 21.05/0.743 parameterized image operator
@@
[…]
@@
[…]
@@
[…]
@@
For Setting5 in Tab. VI, we categorize existing methods into four settings: (a) One-to-One trains on isolated degradations; (b) One-to-Many trains on multiple isolated degradations and tests on them jointly; (c) One-to-Composite trains and tests on both isolated and composite degradations; and (d) Step-by-Step trains on isolated degradations and tests on their combinations step by step. Among the One-to-Composite methods, OneRestore [31], which introduces a novel dataset CDD-11 encompassing both single and composite degradations, achieves a good balance between accuracy and efficiency. In comparison, the AllRestorer model [257] achieves the best performance by adaptive weights and composite scene descriptors. For the Step-by-Step, $\text{CoR}_{\text{OneRestore}}$ demonstrates the advantage of chain of restoration, even only requiring training on $n$ isolated degradations instead of the full $12^{n}-$ 1 combinations. Overall, models in the one-to-composite and step-by-step categories show a certain degree of generalization and robustness, making them worthy of further exploration.
@@
[…]
@@ V-A2 OOD Degradations
V-A3 Model Complexity and Efficienc.
@@ V-B Future Research Directions
Future research will focus on several key directions. First, establishing a robust multi-task learning theory will be essential to effectively address task conflicts and optimize information-sharing mechanisms. Second, reducing reliance on large-scale labeled datasets by exploring semi-supervised and unsupervised learning methods will improve the model’s adaptability in data-scarce situations. Additionally, designing efficient models suitable for edge computing will make AiOIR more feasible for practical applications. Furthermore, investigating more complex real-world degradation scenarios will drive improvements in model performance across various environments. Finally, integrating large multimodal pretrained models and generative priors will enhance the restoration capabilities by leveraging rich multimodal data. These research directions will lay the foundation for the practicality and flexibility of AiOIR models, enabling them to tackle a broader range of real-world challenges.
@@ V-B4 Platform-Aware Model Design and Efficient Approaches
An important direction is developing edge-friendly AiOIR models that balance accuracy and efficiency for deployment on mobile and embedded devices [266, 267]. Techniques like compression, pruning, quantization, and architecture search help reduce resource use with minimal performance loss. In addition to these methods, more radical approaches such as binary neural networks (BNNs) and spiking neural networks (SNNs) offer promising avenues for AiOIR deployment. Recent works show their effectiveness in tasks like super-resolution and deraining [268, 269, 270, 271]. Combined with lightweight transformers, edge-specific distillation, and early-exit strategies, these methods advance practical AiOIR deployment.
@@ V-B9 Extension to High-dimensional Data
AiOIR is increasingly deployed in medical imaging, surveillance, and autonomous driving, yet its ethical implications are seldom addressed. Prior work on machine unlearning [135] shows how to erase the influence of private data in AiOIR, suggesting future research can explore data deletion and privacy-preserving training. Beyond privacy, biased data or model priors can cause fairness issues like uneven performance or hallucinations. Furthermore, misuse risks must be considered, as restored content may be used maliciously in misinformation, surveillance overreach, or unauthorized identity synthesis. These issues highlight the need for responsible development and deployment.
@@ VI Conclusion
In this paper, we present a comprehensive review of recent advancements in all-in-one image restoration, a rapidly emerging domain that integrates multiple types of image degradations into a single framework. Through an in-depth exploration of state-of-the-art models, we emphasize their robust capabilities, diverse architectures, and the rigorous experiments. By contrasting these models with traditional single-task approaches, we underscore the limitations of the latter in addressing real-world complexities, while highlighting the significant gains in efficiency, adaptability, and scalability offered by AiOIR models.
Our comprehensive taxonomy of existing works offers a multi-dimensional perspective, covering structural innovations, key methodologies such as prompt learning, mixture-of-experts, the incorporation of multimodal models, and autonomous agents. We further present an analysis of critical datasets, providing researchers and practitioners with a holistic toolkit to better assess the current landscape of AiOIR. Despite the considerable strides made in recent years, challenges persist. Current models still struggle with handling complex and compounded degradations, lack computational efficiency, and do not generalize well in real-world scenarios. We believe that future research will focus on several key areas: the development of more lightweight and efficient architectures, advancements in semi-supervised learning, and expanding the scope of AiOIR models to accommodate multimodal inputs and video data. Additionally, as the field progresses, innovations in cross-modal learning, real-time processing, and interpretability will likely become central to pushing the boundaries of what AiOIR systems can achieve.
In conclusion, AiOIR represents more than just a technical integration of multiple degradation types—it marks a paradigm shift toward unified, general-purpose vision models that can adapt to real-world complexity with minimal supervision. While significant progress has been made, substantial challenges remain, especially in scaling to real-world conditions, ensuring robustness under compounded degradations, and extending beyond visual-only inputs. We foresee future innovations at the intersection of efficient architecture design, multimodal representation learning, and autonomous adaptation mechanisms. By charting a comprehensive landscape of AiOIR, this review aims not only to inform but also to inspire a new wave of research that pushes the boundaries of what is possible in image restoration and beyond.

### v3

@@ III-B5 Machine Unlearning
Privacy protection is a critical concern in AI, especially as models are deployed in sensitive applications. Machine unlearning [136] offers a way to remove the influence of private data from trained models, enabling them to behave as if such data were never used. Existing methods fall into two categories: exact unlearning [137], which seeks to completely eliminate the impact of specific data points, and approximate unlearning [138], which aims to diminish the influence of such data to a certain degree. Although unlearning has been studied in classification and federated learning, its use in AiOIR remains underexplored.
To fill this gap, Su et al. [139] defined a scenario in which certain types of degradation are treated as private information that must be effectively forgotten by the model. Importantly, this process is designed to preserve the model’s performance on other types of degradations. They introduced instance-wise unlearning, which leveraged adversarial examples in conjunction with gradient ascent methods. This approach not only enhances the AiOIR model’s ability to forget specific data but also maintains its robustness across various image restoration tasks.
@@ III-C Key Improvements of AiOIR
In addition to the network design (Sec. III-A) and learning strategies (Sec. III-B), other key techniques further enhance AiOIR models. We categorize these improvements into four areas: Prompt Learning, Mixture-of-Experts (MoE), Multimodal Model and Agentic IR System, which correspond to Sec. III-C1, Sec. III-C2, Sec. III-C3, Sec. III-C4, respectively. We also highlight additional techniques such as deep unfolding methods, mask image modeling, hypernetworks, etc. At last, we discuss AiOIR for downstream tasks and outline real-world applications. For clarity, representative works for each category are summarized in Fig. 4.
@@ III-C2 Mixture-of-Experts
Recent studies have extended the MoE paradigm to handle weather-related degradations using both language-driven and weather-aware routing mechanisms. For example, Yang et al. [103] generated degradation priors based on textual descriptions of weather conditions, which are then used to guide the adaptive selection of restoration experts. Similarly, WM-MoE [155] introduces a weather-aware routing mechanism to direct image tokens to specialized experts and employs multi-scale experts to handle diverse weather conditions. Both WM-MoE and MoFME [156] enhance IR tasks and downstream tasks like segmentation, while significantly reducing model parameters and inference time. In contrast, MEASNet [86] presents a novel multi-expert adaptive selection mechanism. This mechanism leverages both local and global image features to select the most appropriate expert models, effectively balancing task-specific demands and promoting resource sharing. To further adapt to varying complexities, MoCE-IR [157] introduces complexity experts with adaptive variable-size computational units that use a spring-inspired force mechanism to dynamically allocate experts. Building on similar principles, UniRestorer [158] constructs a multi-granularity degradation set and a multi-granularity MoE restoration model. LoRA-IR [159] leverages a novel mixture of low-rank experts structure, combined with a CLIP-based degradation-guided router. Recently, M2Restore [160] introduces a novel CLIP-guided MoE-based Mamba-CNN framework. Overall, the sparse activation nature of MoE architectures enables more resource-aware AiOIR models with improved inference efficiency and generalizability.
@@ III-C5 Other Methods
In addition to highlighting four key improvements, we also review other methods for AiOIR. Some models benefit from iterative algorithms in deep unfolding networks (DUNs), and integrating network design with pretrained mask image modeling (MIM) also holds significant potential. Meanwhile, hypernetworks with dynamically generated parameters are also highly suitable for AiOIR. The exploration of visual autoregressive modeling (VAR) as a novel paradigm for low-level vision tasks has recently extended to the field of AiOIR. The details are as follows.
Deep Unfolding Methods. Under the framework of optimization theory and deep learning, DUNs bridge interpretability and data-driven flexibility by translating iterative solvers into end-to-end trainable architectures. Zhang et al. [169] proposed a DUN framework that integrates CNNs with model-based priors for task-specific IR. Given a degradation model, the clean image is estimated by minimizing an energy function. The half-quadratic splitting (HQS) algorithm [170] decomposes the optimization into two subproblems: one for the data fidelity term and the other for the prior. These subproblems are alternately solved, with the data term handled via least squares and the prior term approximated using CNNs. Building on this principle, DRM-IR [61] enhances the flexibility of AiO scenarios by introducing a reference-based and task-adaptive modeling paradigm. It establishes an efficient AiOIR framework that jointly models task-adaptive degradation and model-based restoration. Recently, Up-restore [171] employs a prompt-driven unrolling scheme inspired by ADMM and grounded in MAP estimation, effectively combining data-driven learning with unified degradation modeling. Additionally, VLU-Net [153] leverages an automatic gradient estimation strategy based on VLMs, integrating multi-level information to improve feature preservation. These works pave the way for more flexible, efficient, and robust solutions for tackling complex, real-world AiOIR challenges.
@@
  Image Deblurring GoPro [172] 2017 2103 1111 Blurred images at 1280x720 real
@@
Mask Image Modeling. Mask image modeling (MIM) (e.g.,  [200, 148, 201]) is a technique that involves training models to predict masked portions of images based on their surrounding context inspired by mask language modeling [202, 203]. This self-supervised approach reconstructs missing image regions, enhancing its visual representation learning. MIM has demonstrated effectiveness in visual tasks such as classification and segmentation. The MAE [148] framework effectively employs MIM to predict hidden tokens, showcasing impressive performance and generalization capabilities across a range of downstream tasks. Qin et al. [204] are the pioneers in introducing MIM into AiOIR, shifting the focus from degradation‐type discrimination to intrinsic image information recovery. They proposed mask attribute conductance (MAC) to analyze each layer’s contribution to resolving the input integrity gap and rank layers in descending order. DyNet  [205] trains two variants (small and large) concurrently to reconstruct the clean image from masked degraded inputs. Notably, both variants share weights via an intra-network block-repetition scheme, reducing GPU hours by 50%. Most recently, Cat-AIR  [206] introduces a content-aware and task-aware router that dynamically applies self-attention to complex regions and convolution to simpler ones, using learned masks to distinguish hard from easy patches within the feature map.
@@
III-D Applications of AiOIR
Beyond perceptual enhancement, AiOIR methods can also benefit downstream tasks. For example, UniRestore [214] trains a task-aware adapter to align encoder and latent features with various downstream objectives (e.g., classification, semantic segmentation). Similarly, MoFME [156] demonstrates robust performance for both upstream and downstream tasks under adverse weather. Beyond these, AiOIR has promising applications in tasks like depth estimation, remote sensing, medical imaging, and autonomous driving, highlighting the need for further research on practical deployment.
@@ IV Experiments
To facilitate a comprehensive comparison of AiOIR methods, we summarize key datasets and metrics for AiOIR tasks, then compare representative benchmarks under five common experimental setting. This section enables a thorough evaluation of AiOIR methods.
@@ IV-A Datasets and Evaluation Metrics
Single
LPN [226] CVPR’19 3M 20.84/0.828 24.88/0.784 26.47/0.778 24.77/0.748 21.26/0.552 23.64/0.738 Specific
@@
Multiple
MPRNet [229] CVPR’21 16M 28.00/0.958 33.86/0.958 33.27/0.920 30.76/0.871 27.29/0.761 30.63/0.894 General
@@
All-in-One
DL[232] TPAMI’19 2M 26.92/0.931 32.62/0.931 33.05/0.914 30.41/0.861 26.90/0.740 29.98/0.875 parameterized image operator
@@
TextPromptIR [97] $\bigstar\bigoplus$ arXiv’23 124M+110M 31.65/0.978 38.41/0.982 34.17/0.936 31.52/0.893 28.26/0.805 32.80/0.919 textual prompt for AiOIR
InstructIR-3D [98] $\bigstar\bigoplus$ ECCV’24 16M+17M 30.22/0.959 37.98/0.978 34.15/0.933 31.52/0.890 28.30/0.804 32.43/0.913 natural language prompts
@@
DFPIR [247] $\bigstar\bigoplus$ CVPR’25 31M+63M 31.87/0.980 38.65/0.982 34.14/0.935 31.47/0.893 28.25/0.806 32.88/0.919 degradation-aware feature perturbation
@@ IV-B Experiments and Analysis
Single
ADFNet [227] AAAI’23 8M 24.18/0.928 32.97/0.943 31.15/0.882 25.79/0.781 21.15/0.823 27.05/0.871 Specific
@@
Multiple
SwinIR [252] ICCVW’21 1M 21.50/0.891 30.78/0.923 30.59/0.868 24.52/0.773 17.81/0.723 25.04/0.835 General
@@
All-in-One
DL [232] TPAMI’19 2M 20.54/0.826 21.96/0.762 23.09/0.745 19.86/0.672 19.83/0.712 21.05/0.743 parameterized image operator
@@
InstructIR-5D [98] $\bigstar\bigoplus$ ECCV’24 16M+17M 36.84/0.973 27.10/0.956 31.40/0.887 29.40/0.886 23.00/0.836 29.55/0.907 natural language prompts
@@
DFPIR [247] $\bigstar\bigoplus$ CVPR’25 31M+63M 31.64/0.979 37.62/0.978 31.29/0.889 28.82/0.873 23.82/0.843 30.64/0.913 degradation-aware feature perturbation
CyclicPrompt [152] $\bigstar\bigoplus$ arXiv’25 30M+486M 33.02/0.983 37.03/0.989 31.33/0.941 29.42/0.925 21.79/0.837 30.52/0.935 cyclic prompt-driven universal framework
@@
$\mathrm{T}^{3}\text{-}\mathrm{DiffWeather}$ [101] $\bigstar\bigoplus$ 32.37/0.936 31.99/0.937 32.66/0.941 32.34/0.938 69M+25M
@@
For Setting5 in Tab. VI, we categorize existing methods into four settings: (a) One-to-One trains on isolated degradations; (b) One-to-Many trains on multiple isolated degradations and tests on them jointly; (c) One-to-Composite trains and tests on both isolated and composite degradations; and (d) Step-by-Step trains on isolated degradations and tests on their combinations step by step. Among the One-to-Composite methods, OneRestore [32], which introduces a novel dataset CDD-11 encompassing both single and composite degradations, achieves a good balance between accuracy and efficiency. In comparison, the AllRestorer model [261] achieves the best performance by adaptive weights and composite scene descriptors. For the Step-by-Step, $\text{CoR}_{\text{OneRestore}}$ demonstrates the advantage of chain of restoration, even only requiring training on $n$ isolated degradations instead of the full $2^{n}-1$ combinations. Overall, models in the one-to-composite and step-by-step categories show a certain degree of generalization and robustness, making them worthy of further exploration.
@@
AllRestorer [261] $\bigstar\bigoplus$ arXiv’24 33.72 0.944 12M+88M
$\text{DA-CLIP}_{\text{NAFNet}}$ $\bigstar$ ICLR’24 26.56 0.860 86M+125M
@@ V-A2 OOD Degradations
V-A3 Model Complexity and Efficiency
@@ V-B Future Research Directions
Future research will pursue several directions: developing a robust multi-task learning theory to resolve task conflicts and enhance information sharing; exploring semi‑ and unsupervised methods to reduce dependence on large labeled datasets; designing efficient, edge‑compatible models for practical deployment; examining complex real‑world degradation scenarios to boost performance; and integrating large multimodal pretrained models with generative priors. Together, these efforts will underpin more practical, flexible AiOIR models capable of addressing diverse real‑world challenges.
@@ V-B4 Platform-Aware Model Design and Efficient Approaches
An important direction is developing edge-friendly AiOIR models that balance accuracy and efficiency for deployment on mobile and embedded devices [270, 271]. Techniques like compression, pruning, quantization, and architecture search help reduce resource use with minimal performance loss. In addition to these methods, more radical approaches such as binary neural networks (BNNs) and spiking neural networks (SNNs) offer promising avenues for deployment. Recent works show their effectiveness in tasks like super-resolution and deraining [272, 273, 274, 275]. Combined with lightweight transformers, edge-specific distillation, and early-exit strategies, these methods advance practical AiOIR deployment.
@@ V-B9 Extension to High-dimensional Data
AiOIR is increasingly deployed in medical imaging, surveillance, and autonomous driving, yet its ethical implications are rarely addressed. Prior work on machine unlearning [139] shows how to erase the influence of private data in AiOIR, suggesting future research can explore data deletion and privacy-preserving training. Beyond privacy, biased data or model priors can cause fairness issues like uneven performance or hallucinations. Furthermore, misuse risks must be considered, as restored content may be used maliciously in misinformation, surveillance overreach, or unauthorized identity synthesis. These issues highlight the need for responsible development and deployment.
@@ VI Conclusion
In this paper, we present a comprehensive review of recent advancements in all-in-one image restoration (AiOIR), a rapidly emerging domain that integrates multiple types of image degradations into a single framework. Through an in-depth exploration of state-of-the-art models, we emphasize their robust capabilities, diverse architectures, and the rigorous experiments. By contrasting these models with traditional single-task approaches, we underscore the limitations of the latter in addressing real-world complexities, while highlighting the significant gains in efficiency, adaptability, and scalability offered by AiOIR models.
Our proposed taxonomy offers a multi-faceted perspective on this rapidly evolving field, covering network innovations, key methodologies such as prompt learning, mixture-of-experts, the incorporation of multimodal models, and autonomous agents. We further present an analysis of critical datasets, providing researchers and practitioners with a holistic toolkit to better assess the current landscape of AiOIR. Despite notable advancements, persistent challenges remain: current models still struggle with handling complex and compounded degradations, suffer from computational inefficiencies, and do not generalize well in real-world scenarios. Looking ahead, we anticipate that the trajectory of AiOIR research will be shaped by several converging trends: the development of more lightweight and efficient architectures, advancements in semi-supervised learning, and expanding the scope of AiOIR models to accommodate multimodal inputs and video data. Furthermore, breakthroughs in real-time processing, cross-modal reasoning, and model interpretability are poised to play a pivotal role in unlocking the full potential of unified restoration frameworks.
In conclusion, AiOIR is not merely a technical amalgamation of multiple restoration tasks—it represents a foundational rethinking of how visual systems can perceive, process, and reconstruct information in complex environments. By charting the contours of this evolving landscape, our review seeks not only to synthesize current knowledge but also to catalyze future innovations at the intersection of efficient computation, multimodal understanding, and autonomous adaptation. As AiOIR continues to mature, it holds the promise of becoming a cornerstone technology for general-purpose visual intelligence in the wild.